ganeshie8 (ganeshie8):

show that \[\large \sum_{n=1}^{\infty}\arctan\frac{1}{2n^2}=\frac{\pi}{4}\]

ganeshie8 (ganeshie8):

spoiler hint : \(\huge \color{white}{🔭}\)

OpenStudy (freckles):

\[\tan(\frac{\pi}{4})=1 \\ \tan(\sum \arctan(\frac{1}{2n^2}))=1 \\ \tan(\arctan(\frac{1}{2})+\arctan(\frac{1}{8})+\arctan(\frac{1}{18})+\cdots) \text{ w.t.s this is }1 \] but that might be harder to show not sure yet

OpenStudy (freckles):

\[\tan(\arctan(\frac{1}{2})+s_1)=\frac{\tan(\arctan(\frac{1}{2}))+\tan(s_1)}{1-\tan(\arctan(\frac{1}{2})) \tan(s_1)}=\frac{\frac{1}{2}+\tan(s_1)}{1-\frac{1}{2} \tan(s_1)}\\ =\frac{1+2 \tan(s_1)}{2-\tan(s_1)}\]

OpenStudy (freckles):

where \[s_1=\arctan(\frac{1}{8})+\arctan(\frac{1}{18})+ \cdots \]

OpenStudy (freckles):

yep seems like a dead end for me

ganeshie8 (ganeshie8):

thats prety clever! i think that might actually telescope... im still trying on paper..

ganeshie8 (ganeshie8):

\[\arctan\frac{1}{2n^2} \\~\\= \arctan\frac{2}{(2n)^2} \\~\\= \arctan\frac{2}{1+(2n)^2-1}\\~\\=\arctan\dfrac{(2n+1)-(2n-1)}{1+(2n+1)(2n-1)}\]

ganeshie8 (ganeshie8):

\[\arctan(\dfrac{x-y}{1+xy}) = \arctan(x)-\arctan(y)\] let \(x=(2n+1)\) and \(y=2n-1\)

OpenStudy (anonymous):

this is nice too, classic

ganeshie8 (ganeshie8):

im biased to telescoping xD but there is another way to work this using complex numbers which is pretty neat too the basic idea is to interpret the "multiplication of complex numbers" geometrically by "adding up the angles"

ganeshie8 (ganeshie8):

\[\arg(z_1z_2) = \arg(z_1) + \arg(z_2)\]

OpenStudy (astrophysics):

That sounds fascinating, have you written it down, can you just post a pic of it, then you don't have to write it out

OpenStudy (astrophysics):

on here

ganeshie8 (ganeshie8):

it is easy, just need to see it one time that the angles add up when we multiply a complex number by another complex number :